A hole of cross-sectional area 100cm² is cut in a piece of steel at 20⁰c.

What is the area of the hole if the steel is heated from 20⁰c to 100⁰c?

The coefficient of linear expression of steel is 1.10 * 10–5⁰c


1
Expert's answer
2021-10-22T08:29:09-0400

"S=S_0(1+2\alpha \Delta t),"

"S=100\cdot (1+2\cdot 1.1\cdot 10^{-5}\cdot (100-20))=100.176~cm^2."